Determine the viscosity of a liquid having kinematic viscosity 6 stokes and specific gravity 2.0
1
Expert's answer
2020-11-09T19:04:10-0500

Specific gravity


"S(for\ liquids)=\dfrac{\rho_L}{\rho_{H_2O}}"

Density of water "\rho_{H_2O}=1000kg/m^3"

Given "S=2.0, \nu_L=6St=6\times10^{-4}m^2/s"

Use the relation


"\nu_L=\dfrac{\mu_L}{\rho_L}"

"\mu_L=\nu_L\cdot\rho_L=\nu_L\cdot S\cdot\rho_{H_2O}"

"\mu_L=6\times10^{-4}m^2/s\cdot2\cdot1000kg/m^3="

"=1.2Ns/m^2=12\text{ poise}"